Six Measures of Solvency or Profitability

The following data were taken from the financial statements of Gates Inc. for the current fiscal year.

Property, plant, and equipment (net) $1,715,400
Liabilities:
Current liabilities $190,000
Note payable, 6%, due in 15 years 953,000
Total liabilities $1,143,000
Stockholders' equity:
Preferred $4 stock, $100 par (no change during year) $1,714,500
Common stock, $10 par (no change during year) 1,714,500
Retained earnings:
Balance, beginning of year $1,828,000
Net income 567,000 $2,395,000
Preferred dividends $68,580
Common dividends 40,420 109,000
Balance, end of year 2,286,000
Total stockholders' equity $5,715,000
Sales $32,095,200
Interest expense $57,180

Assuming that total assets were $6,515,000 at the beginning of the current fiscal year, determine the following. When required, round to one decimal place.

a. Ratio of fixed assets to long-term liabilities
b. Ratio of liabilities to stockholders' equity
c. Asset turnover
d. Return on total assets %
e. Return on stockholders’ equity %
f. Return on common stockholders' equity %
Answer #1
a. Ratio of fixed assets to long term liabilties = Fixed assets / Long term liabilities = 1715400 / 953000 1.8
b. Ratio of liabilities to stockholders' equity = Liabilities / Stockholders' equity = 1143000 / 5715000 0.2
c. Ending total assets = Ending total liabilities + Ending stockholders' equity = 1143000 + 5715000 6858000
Average total assets = ( Beginning total assets + Ending total assets ) / 2 = ( 6515000 + 6858000 ) / 2 6686500
Asset turnover = Sales / Average total assets 4 4.8
d. Return on total assets = Net income / Average Total assets = 567000 / 6686500 8.5%
e. Beginning stockholders' equity = Preferred stock + Common stock + Beginning retained earnings = 1714500 + 1714500 + 1828000 5257000
Average stockholders' equity = ( Beginning stockholders' equity + Ending stockholders' equity ) / = ( 5257000 + 5715000 ) / 2 5486000
Return on stockholders' equity = Net income / Average Stockholders' equity = 567000 / 5486000 10.3%
f. Beginning Ending
Stockholders' equity 5257000 5715000
(-) Preferred stock 1714500 1714500
Common stockholders' equity 3542500 4000500
Average common stockholders' equity = ( Beginning common stockholders' equity + Ending common stockholders' equity ) / 2 = ( 3542500 + 4000500 ) / 2 3771500
Return on common stockholders' equity = ( Net income - Preferred dividends ) / Average common stockholders' equity = ( 567000 - 68580 ) / 3771500 13.2%